# Designing LLM Chains by Adapting Techniques from Crowdsourcing Workflows

### v3

@@ 4.1.2. Accuracy vs creativity
One element of output quality of particular importance is if the workflow should support the accuracy or creativity of outputs. Many tasks’ definitions of output quality have elements of both accuracy and creativity.
@@
Accuracy. Accuracy-focused tasks have one correct answer. Some crowdsourcing papers in this space include TurKit (Little et al., 2010b) and Second Opinion (Mohanty et al., 2019). Examples of these tasks include math problems (Zhou et al., 2022b; Liang et al., 2023; Du et al., 2023; Wu et al., 2023b; Madaan et al., 2023b), blurry text recognition (Little et al., 2010b, a), and matching images with the same content (Mohanty et al., 2019; Kobayashi et al., 2018). For these tasks, workflows can improve performance on benchmarks (Schick et al., 2022; Guo et al., 2019; Zhang et al., 2022; Cobbe et al., 2021; Li et al., 2023; Wu et al., 2023b; Eisenstein et al., 2022; McDonnell et al., 2016; Kobayashi et al., 2018), enable new tasks (Guo et al., 2017; Huffaker et al., 2020; Gouravajhala et al., 2018; Little et al., 2010b; Huang et al., 2023c), classify examples for which machine learning approaches struggle (Mohanty et al., 2019; Bernstein et al., 2010; Cheng and Bernstein, 2015; Alshaibani et al., 2020), improve instructional clarity (Bragg et al., 2018), and reduce hallucinations (Nair et al., 2023a; Mirowski et al., 2023; Creswell et al., 2022; Creswell and Shanahan, 2022; Nair et al., 2023b; Wu et al., 2023a; Parameswaran et al., 2023).
@@
Creativity. Creative tasks have the widest degree of output freedom. Representative crowdsourcing papers for creative objectives include Flash Teams (Retelny et al., 2014), VisiBlends (Chilton et al., 2019), and Mechanical Novel (Kim et al., 2017). These papers tackle creative writing (André et al., 2014; Kim et al., 2017; Bernstein et al., 2010; Yao et al., 2023; Yang et al., 2022; Mirowski et al., 2023; Zhang et al., 2023a), idea generation (Little et al., 2010a; Kim et al., 2023b; Wu et al., 2023c), and prototyping (Lee et al., 2017; Kulkarni et al., 2014; Valentine et al., 2017; Retelny et al., 2014; Park et al., 2022). Prior work finds that on creative tasks, workflows can help achieve higher quality (André et al., 2014; Kim et al., 2017; Yao et al., 2023; Yang et al., 2022; Mirowski et al., 2023; Kim et al., 2023b; Lee et al., 2017; Park et al., 2022) and less constrained outputs (Liu et al., 2018; Hahn et al., 2016; Lee et al., 2017; Drapeau et al., 2016; Wang et al., 2023; Park et al., 2022). Workflows may enable novice crowdworkers (Pietrowicz et al., 2013) or users (Wang et al., 2023; Lin et al., 2014a; André et al., 2013; Wu et al., 2023c; Valentine et al., 2017; Chilton et al., 2019; Kim and Monroy-Hernandez, 2016) to perform a creative task with an otherwise high barrier of entry. Important qualities for creative tasks include coherence (Yao et al., 2023; Yang et al., 2022; Kim et al., 2017) and aligning to the user’s creative vision (Salehi et al., 2017). On creative tasks, a workflow is often better suited for creating and lengthening the output, rather than proofreading (Kim et al., 2017; Huang et al., 2017; Little et al., 2010b; Mirowski et al., 2023), so these outputs are sometimes seen as a starting point for editing, rather than a finished product (Kim and Monroy-Hernandez, 2016; Park et al., 2022; Long et al., 2023; Mirowski et al., 2023; Nebeling et al., 2016).
@@
Both accuracy and creativity. Many tasks require some degree of accuracy and some degree of creativity. In some tasks, there is not one correct answer, but there are restrictions on what outputs are correct. Crowdsourcing papers for these tasks include Cascade (Chilton et al., 2013), Flock (Cheng and Bernstein, 2015), Revolt (Chang et al., 2017), and Storia (Kim and Monroy-Hernandez, 2016). Example tasks include categorizing data (categories must be accurate, but there are multiple options and unclear edge cases) (Bragg et al., 2013; Chilton et al., 2013; André et al., 2013; Chilton et al., 2014; Cheng and Bernstein, 2015; Chang et al., 2017), summarizing text (summaries must stay faithful, but have degrees of freedom in how to write the summary) (Kim and Monroy-Hernandez, 2016; Pavel et al., 2014; Bernstein et al., 2010; Nair et al., 2023a, b), and coding (programs must work, but there are multiple approaches to problem solving) (LaToza et al., 2014; Kulkarni et al., 2014; Wu et al., 2023a; Huang et al., 2023b, a). The goals of workflows for these tasks are to align with expert outputs (Kim et al., 2014; Liu et al., 2018; Chilton et al., 2013; Salehi et al., 2017; Park et al., 2022; Kim et al., 2023b; Fleming et al., 2023) or to enable users to choose among multiple correct possible output options. Using workflows for these tasks can result in an output that is human-understandable (Cheng and Bernstein, 2015), that better represents what would be important for people to know (Salisbury et al., 2017), and that surfaces the uncertainty among the different possible correct options (Chang et al., 2017).
@@ 4.1.3. Task specificity vs generality
Some workflows are highly specialized to the specifications of the task (e.g., visual blends (Wang et al., 2023; Chilton et al., 2019), masking faces in images (Alshaibani et al., 2020; Kaur et al., 2017), and programming (Huang et al., 2023b; LaToza et al., 2014)). Others generalize to a variety of tasks (Creswell et al., 2022; Creswell and Shanahan, 2022; Nair et al., 2023b; Kulkarni et al., 2012; Wu et al., 2023a). For example, AutoGen (Wu et al., 2023a) uses the same conversation-based workflow for coding as well as chess games, math problems, and retrieval-based question answering. Task-specific workflows support fewer outcome types but can incorporate more tailored support than workflows that generalize to many tasks.
Title: Tactics, the building blocks of workflows. One row each for workers, subtasks, and architectures, with the category and (a short description). The workers row contains: Crowdworkers (people on an online platform), LLMs (generative large language models), user (the person deploying the workflow). The subtasks row contains: generate (create content), evaluate (judge content), transform (update content), focus (select content) and partition (split tasks). The architecture row contains: sequential (call LLMs in sequence), branching (diverge on paths), redundant (multiple workers repeat subtask), dynamic (Initialize parts of the workflow at runtime), and communicative (facilitate interaction among workers) Each subtask category is connected with a generic example instruction. Generate: ”Write a caption”. Evaluate: ”Is this answer grammatical?”. Transform: ”Improve the given idea”. Focus: ”Select regions for future work:. Partition: ”Divide the task into subtasks”.
@@ 4.2. Tactics in the design space
Designers use workflow tactics to implement workflows. Tactics need to consider the workers, the subtasks the workers complete, and the connections of subtasks into workflow architectures (Figure 3).
4.2.1. Workers
Workers may be crowdworkers, LLMs, or the user.
@@ 4.2.1. Workers
Crowdworkers. These workers are usually recruited from crowdworking platforms like Amazon Mechanical Turk or UpWork (formerly known as oDesk). Platforms allow for requests of worker qualifications, although these credentials are sometimes hard to validate (Ahmad et al., 2011; Zaidan and Callison-Burch, 2011; Valentine et al., 2017; Kulkarni et al., 2014). Methods like retainer models can reduce the latency of accessing crowdworkers on these platforms (Bernstein et al., 2011). Workers can also be recruited from the public, including community members (Mahyar et al., 2018), conference committees (Chilton et al., 2014; André et al., 2013), and online deaf communities (Bragg et al., 2021).
Crowdworkers have limitations that overlap with LLM limitations. Prior work suggests that approximately 30% of outcomes from Amazon Mechanical Turk are low quality (Bernstein et al., 2010), and may include spam or off-topic responses (Butler et al., 2017; Lasecki et al., 2013; Alshaibani et al., 2020; Huang et al., 2018). Crowdworkers also struggle with following long and complex instructions, as people can overly focus on the beginning or end of a given text (Hahn et al., 2016; Chilton et al., 2013; Lasecki et al., 2015; Chen et al., 2019). Crowdworkers can be biased; for example, in one study, workers performed significantly better in identifying photos of their own race compared to other races (Mohanty et al., 2019).
Crowdworkers have unique limitations. Anchoring bias can lead to less creativity in generation or divergent thinking (Wu et al., 2022b, 2023c). Since workers are paid when they complete a task, there is incentive to work quickly even if that means satisficing (Bernstein et al., 2010; Parameswaran et al., 2023; Hata et al., 2017). Additionally, since rejected work can hurt access to future jobs on a platform (Kim et al., 2017), workers may overcompensate to a detrimental level in an attempt to make their efforts clear (Kim et al., 2014; Bernstein et al., 2010). Crowdworkers may fatigue (Alshaibani et al., 2020; Dai et al., 2013), and they may lack task-relevant knowledge and delete jargon (Nebeling et al., 2016; Bernstein et al., 2010). On synchronous tasks, other limitations become prevalent. Social blocking can make workers so cautious to change others’ work that not enough work gets done (Lasecki et al., 2015; Hahn et al., 2016), and territoriality can motivate edit wars in which workers’ attempts to establish their own work over others degrade the final product quality (Valentine et al., 2017; André et al., 2014). Crowdworkers also may be affected by diffusion of responsibility and social loafing in which there is incentive to wait for others to do tasks (Lasecki et al., 2015; André et al., 2014).
Crowdworkers have their comparative strengths as well, including the ability to perform high-level structural evaluation and to quickly identify outliers (Huang et al., 2015; Luther et al., 2015). People are also currently better at selectively choosing among subtasks than LLMs (Wu et al., 2023c).
@@
LLMs. There are many possible LLMs of many sizes and expenses (Bursztyn et al., 2022). Models have differing strengths and respond differently to prompts (Mirowski et al., 2023; Wu et al., 2023c; Parameswaran et al., 2023; Liang et al., 2023). LLM outputs can be adjusted through parameter changes (e.g., temperature) (Park et al., 2022; Parameswaran et al., 2023), fine-tuning (Creswell et al., 2022; Creswell and Shanahan, 2022; Cobbe et al., 2021; Li et al., 2023; Bursztyn et al., 2022), and expressing roles through prompts (Park et al., 2022; Du et al., 2023).
@@
Some LLM limitations overlap with those of crowdworkers. LLMs can provide erroneous outputs: they are known to confidently output incorrect answers (Parameswaran et al., 2023; Wang et al., 2023; Du et al., 2023) or hallucinations (Nair et al., 2023a; Mirowski et al., 2023; Creswell et al., 2022; Creswell and Shanahan, 2022; Nair et al., 2023b; Wu et al., 2023a; Parameswaran et al., 2023) and for struggling with self-consistency and logical reasoning (Creswell et al., 2022; Creswell and Shanahan, 2022; Yao et al., 2023; Parameswaran et al., 2023; Liang et al., 2023; Xie et al., 2023). LLMs can ignore large fractions of long prompts or focus mostly on recent parts of the given context (Parameswaran et al., 2023; Du et al., 2023; Wu et al., 2023c; Madaan et al., 2023b). LLMs also output biased results, a widespread concern (Park et al., 2022; Wu et al., 2023a; Huang et al., 2023a; Long et al., 2023). For example, work on screenplay generation found stereotypical outputs, with multiple instances of gender biases (Mirowski et al., 2023). Furthermore, LLM training on human data increases the risk of accidental plagiarism (Mirowski et al., 2023).
@@
LLMs also have unique limitations. On creative tasks, LLMs sometimes produce outputs that are perceived as robotic (Long et al., 2023) and that can lack the understanding, nuance, and subtext that people can provide (Mirowski et al., 2023). Furthermore, models can get stuck in generative loops, degenerating text quality (Parameswaran et al., 2023). LLMs also hallucinate more than crowds (Nair et al., 2023a; Mirowski et al., 2023; Creswell et al., 2022; Creswell and Shanahan, 2022; Nair et al., 2023b; Wu et al., 2023a; Parameswaran et al., 2023) and may struggle to change their outputs through self-reflection (Liang et al., 2023).
@@
LLMs have comparative strengths as well. They may be better at divergent thinking (Wu et al., 2023a, c; Liang et al., 2023) and comparison-based instructions (Wu et al., 2022b, 2023c) than crowdworkers, who are affected more by anchoring bias. LLMs may also be better than crowdworkers at tasks requiring recollection of large amounts of data (Parameswaran et al., 2023).
@@
User. Finally, users of the workflow can also be involved to revise or guide the process. Users can be novices using the workflow to adapt to a steep learning curve (Kulkarni et al., 2014; Pietrowicz et al., 2013; Lin et al., 2014a; Chilton et al., 2019; Wang et al., 2023), or experts using the workflow as a support tool (Mirowski et al., 2023; Nebeling et al., 2016; Mohanty et al., 2019). Involving users in completing subtasks grants them greater control at the expense of higher effort (Stretcu et al., 2023). However, even with user involvement, prior work finds mixed results on how well tools retain the user’s sense of agency and ownership (Kim et al., 2023b; Mirowski et al., 2023).

### v4

@@ 4.1.2. Objectivity vs subjectivity
One element of output quality of particular importance is if the workflow should support objective or subjective outputs. Many tasks’ definitions of output quality have elements of both objectivity and subjectivity.
@@
Objectivity. Objectivity-focused tasks have one correct answer. Some crowdsourcing papers in this space include TurKit [C(Little et al., 2010b)] and Second Opinion [C(Mohanty et al., 2019)]. Examples of these tasks include math problems [L(Zhou et al., 2022b; Du et al., 2023; Wu et al., 2024; Liang et al., 2023; Madaan et al., 2024)], blurry text recognition [C(Little et al., 2010b, a)], and matching images with the same content [C(Kobayashi et al., 2018; Mohanty et al., 2019)]. For these tasks, workflows can improve performance on benchmarks [C(Guo et al., 2019; McDonnell et al., 2016; Kobayashi et al., 2018) L(Li et al., 2023; Eisenstein et al., 2022; Wu et al., 2024; Zhang et al., 2022; Schick et al., 2022; Cobbe et al., 2021)], enable new tasks [C(Little et al., 2010b; Gouravajhala et al., 2018; Huffaker et al., 2020; Guo et al., 2017) L(Huang et al., 2023c)], classify examples for which machine learning approaches struggle [C(Bernstein et al., 2010; Alshaibani et al., 2020; Cheng and Bernstein, 2015; Mohanty et al., 2019)], and reduce hallucinations [L(Wu et al., 2023a; Creswell et al., 2022; Nair et al., 2023a; Creswell and Shanahan, 2022; Parameswaran et al., 2023; Nair et al., 2023b; Mirowski et al., 2023)].
@@
Subjectivity. Subjective tasks have the widest degree of output freedom. Representative crowdsourcing papers with subjective outputs include Flash Teams [C(Retelny et al., 2014)], VisiBlends [C(Chilton et al., 2019)], and Mechanical Novel [C(Kim et al., 2017)]. These papers tackle creative writing [C(Kim et al., 2017; André et al., 2014; Bernstein et al., 2010) L(Yao et al., 2024; Yang et al., 2022; Mirowski et al., 2023; Zhang et al., 2023a)], idea generation [C(Little et al., 2010a) L(Wu et al., 2023b; Kim et al., 2023b)], and prototyping [C(Kulkarni et al., 2014; Valentine et al., 2017; Lee et al., 2017; Retelny et al., 2014) L(Park et al., 2022)]. Prior work finds that on subjective tasks, workflows can help achieve higher quality [C(Kim et al., 2017; André et al., 2014; Lee et al., 2017) L(Yao et al., 2024; Yang et al., 2022; Kim et al., 2023b; Mirowski et al., 2023; Park et al., 2022)] and less constrained outputs [C(Liu et al., 2018; Hahn et al., 2016; Drapeau et al., 2016; Lee et al., 2017) L(Park et al., 2022; Wang et al., 2023)]. Workflows may enable novice crowdworkers [C(Pietrowicz et al., 2013)] or users [C(Chilton et al., 2019; Valentine et al., 2017; André et al., 2013; Kim and Monroy-Hernandez, 2016; Lin et al., 2014a) L(Wu et al., 2023b; Wang et al., 2023)] to perform a subjective task with an otherwise high barrier of entry. Important qualities for subjective tasks include coherence [C(Kim et al., 2017) L(Yang et al., 2022; Yao et al., 2024)] and aligning to the user’s creative vision [C(Salehi et al., 2017)]. On subjective tasks, a workflow is often better suited for creating and lengthening the output, rather than proofreading [C(Huang et al., 2017; Kim et al., 2017; Little et al., 2010b) L(Mirowski et al., 2023)], so these outputs are sometimes seen as a starting point for editing, rather than a finished product [C(Nebeling et al., 2016; Kim and Monroy-Hernandez, 2016) L(Mirowski et al., 2023; Long et al., 2023; Park et al., 2022)].
@@
Both objectivity and subjectivity. Many tasks require some degree of objectivity and some degree of subjectivity. In some tasks, there is not one correct answer, but there are restrictions on what outputs are correct. Crowdsourcing papers for these tasks include Cascade [C(Chilton et al., 2013)], Flock [C(Cheng and Bernstein, 2015)], Revolt [C(Chang et al., 2017)], and Storia [C(Kim and Monroy-Hernandez, 2016)]. Example tasks include categorizing data (categories must be accurate, but there are multiple options and unclear edge cases) [C(Chilton et al., 2014; André et al., 2013; Bragg et al., 2013; Chilton et al., 2013; Cheng and Bernstein, 2015; Chang et al., 2017)], summarizing text (summaries must stay faithful, but have degrees of freedom in how to write the summary) [C(Kim and Monroy-Hernandez, 2016; Pavel et al., 2014; Bernstein et al., 2010) L(Nair et al., 2023a, b)], and coding (programs must work, but there are multiple approaches to problem solving) [C(Kulkarni et al., 2014; LaToza et al., 2014) L(Wu et al., 2023a; Huang et al., 2023b, a)]. The goals of workflows for these tasks are to align with expert outputs [C(Liu et al., 2018; Kim et al., 2014; Chilton et al., 2013; Salehi et al., 2017) L(Kim et al., 2023b; Park et al., 2022; Fleming et al., 2023)] or to enable users to choose among multiple correct possible output options. Using workflows for these tasks can result in an output that is human-understandable [C(Cheng and Bernstein, 2015)], that better represents what would be important for people to know [C(Salisbury et al., 2017)], and that surfaces the uncertainty among the different possible correct options [C(Chang et al., 2017)].
@@ 4.1.3. Task specificity vs generality
Some workflows are highly specialized to the specifications of the task (e.g., visual blends [C(Chilton et al., 2019) L(Wang et al., 2023)], masking faces in images [C(Alshaibani et al., 2020; Kaur et al., 2017)], and programming [C(LaToza et al., 2014) L(Huang et al., 2023b)]). Others generalize to a variety of tasks [C(Kulkarni et al., 2012) L(Wu et al., 2023a; Creswell et al., 2022; Creswell and Shanahan, 2022; Nair et al., 2023b)]. For example, AutoGen [L(Wu et al., 2023a)] uses the same conversation-based workflow for coding as well as chess games, math problems, and retrieval-based question answering. Task-specific workflows support fewer outcome types but can incorporate more tailored support than workflows that generalize to many tasks.
@@ 4.2. Tactics in the design space
Designers use workflow tactics to implement workflows. Tactics need to consider the actors, the subtasks the actors complete, and the connections of subtasks into workflow architectures (Figure 3).
4.2.1. Actors
Actors may be crowdworkers, LLMs, or the user.
@@ 4.2.1. Actors
Crowdworkers. These actors are usually recruited from crowdworking platforms like Amazon Mechanical Turk or UpWork (formerly known as oDesk). Platforms allow for requests of crowdworker qualifications, although these credentials are sometimes hard to validate [C(Kulkarni et al., 2014; Zaidan and Callison-Burch, 2011; Valentine et al., 2017; Ahmad et al., 2011)]. Methods like retainer models can reduce the latency of accessing crowdworkers on these platforms [C(Bernstein et al., 2011)]. Workers can also be recruited from the public, including community members [C(Mahyar et al., 2018)], conference committees [C(Chilton et al., 2014; André et al., 2013)], and online deaf communities [C(Bragg et al., 2021)].
@@
LLMs. There are LLMs of many sizes and expenses [L(Bursztyn et al., 2022)]. Models have varying strengths and respond differently to prompts [L(Parameswaran et al., 2023; Wu et al., 2023b; Mirowski et al., 2023; Liang et al., 2023)]. LLM outputs can be adjusted through parameter changes (e.g., temperature) [L(Parameswaran et al., 2023; Park et al., 2022)], fine-tuning [L(Li et al., 2023; Creswell et al., 2022; Bursztyn et al., 2022; Creswell and Shanahan, 2022; Cobbe et al., 2021)], and expressing roles through prompts [L(Du et al., 2023; Park et al., 2022)]. Recently, designers often substitute specialized or distilled models, such as embeddings, for LLM functionality at reduced cost [L(Mirowski et al., 2023; Li et al., 2023)].
@@
Comparing crowdworkers and LLMs. Some limitations of crowdworkers and LLMs overlap. First, both crowdworkers and LLMs can output low quality responses. Prior work finds that outcomes from crowdsourcing platforms include low-quality, spam, and off-topic responses [C(Lasecki et al., 2013; Alshaibani et al., 2020; Huang et al., 2018; Butler et al., 2017; Bernstein et al., 2010)]. LLMs can provide erroneous outputs: they are known to confidently output incorrect answers [L(Parameswaran et al., 2023; Du et al., 2023; Wang et al., 2023)] and to struggle with self-consistency and logical reasoning [L(Creswell et al., 2022; Yao et al., 2024; Creswell and Shanahan, 2022; Parameswaran et al., 2023; Liang et al., 2023; Xie et al., 2023)]. Second, long and complex instructions can lower performance for both crowdworkers and LLMs. Crowdworkers struggle with following long and complex instructions, as people can overly focus on the beginning or end of a given text [C(Hahn et al., 2016; Chilton et al., 2013; Chen et al., 2019; Lasecki et al., 2015)]. LLMs can ignore large fractions of long prompts or focus mostly on recent parts of the given context [L(Parameswaran et al., 2023; Du et al., 2023; Wu et al., 2023b; Madaan et al., 2024)]. Third, both crowdworkers and LLMs can give biased outputs. Crowdworkers can be biased; for example, in one study, crowdworkers performed significantly better in identifying photos of their own race compared to other races [C(Mohanty et al., 2019)]. LLMs also output biased results, a widespread concern [L(Wu et al., 2023a; Huang et al., 2023a; Long et al., 2023; Park et al., 2022)]. For example, work on screenplay generation found stereotypical outputs, with multiple instances of gender biases [L(Mirowski et al., 2023)].
Crowdworkers have some unique limitations. Since crowdworkers are paid when they complete a task, there is an incentive to work quickly even if that means satisficing [C(Bernstein et al., 2010) L(Parameswaran et al., 2023)]. Additionally, since rejected work can hurt access to future jobs on a platform [C(Kim et al., 2017)], crowdworkers may overcompensate to a detrimental level in an attempt to make their efforts clear [C(Bernstein et al., 2010; Kim et al., 2014)]. Crowdworkers may fatigue [C(Alshaibani et al., 2020; Dai et al., 2013)], and they may lack task-relevant knowledge and delete jargon [C(Nebeling et al., 2016; Bernstein et al., 2010)]. On synchronous tasks, other limitations become prevalent. Social blocking can make crowdworkers so cautious to change others’ work that not enough work gets done [C(Hahn et al., 2016; Lasecki et al., 2015)], and territoriality can motivate edit wars in which crowdworkers’ attempts to establish their own work over others degrade the final product quality [C(André et al., 2014; Valentine et al., 2017)]. Crowdworkers also may be affected by diffusion of responsibility and social loafing in which there is incentive to wait for others to do tasks [C(André et al., 2014; Lasecki et al., 2015)].
@@
LLMs also have unique limitations. On creative tasks, LLMs sometimes produce outputs that are perceived as robotic [L(Long et al., 2023)] and that can lack the understanding, nuance, and subtext that people can provide [L(Mirowski et al., 2023)]. LLM training on human data increases the risk of accidental plagiarism [L(Mirowski et al., 2023)]. Furthermore, models can get stuck in generative loops, degenerating text quality [L(Parameswaran et al., 2023)]. LLMs also hallucinate more than crowds [L(Wu et al., 2023a; Creswell et al., 2022; Nair et al., 2023a; Creswell and Shanahan, 2022; Parameswaran et al., 2023; Nair et al., 2023b; Mirowski et al., 2023)] and may struggle to change outputs through self-reflection [L(Liang et al., 2023)].
@@
Crowdworkers and LLMs have their comparative strengths as well. Crowdworkers excel at performing high-level structural evaluation and quickly identifying outliers [C(Luther et al., 2015; Huang et al., 2015)]. People are also currently better at selectively choosing among subtasks than LLMs [L(Wu et al., 2023b)]. LLMs may be better at divergent thinking [L(Wu et al., 2023a, b; Liang et al., 2023)] and comparison-based instructions [L(Wu et al., 2023a, b)] than crowdworkers, who are affected more by anchoring bias. LLMs may also be better than crowdworkers at tasks requiring recollection of large amounts of data [L(Parameswaran et al., 2023)].
@@
User. Finally, users of the workflow can also be involved in revising or guiding the process. Users can be novices using the workflow to adapt to a steep learning curve [C(Kulkarni et al., 2014; Chilton et al., 2019; Pietrowicz et al., 2013; Lin et al., 2014a) L(Wang et al., 2023)], or experts using the workflow as a support tool [C(Nebeling et al., 2016; Mohanty et al., 2019) L(Mirowski et al., 2023)]. However, even with user involvement, prior work finds mixed results on how well tools retain the user’s sense of agency and ownership [L(Mirowski et al., 2023; Kim et al., 2023b)].